OpenStudy (anonymous):

9^4sqrt7 - 8^3sqrt7 + 4^3sqrt7 + 3^4sqrt7

OpenStudy (driftracer305):

is it \[9^4.............or ............9 \sqrt[4]{7}\]

OpenStudy (anonymous):

the second one @driftracer305

OpenStudy (driftracer305):

ahh so simple:-- \[12\sqrt[4]{7} - 4\sqrt[3]{7}\]

OpenStudy (anonymous):

this one was HARD what are you saying "easy" ahahaha

OpenStudy (driftracer305):

lol @chariety101 ......it's always like that with me...... i can solve harder ones faster than easy ones..........guess m just used to my higher lvl math thinking.....

jimthompson5910 (jim_thompson5910):

Let \[\large x = \sqrt[4]{7}\] \[\large y = \sqrt[3]{7}\] ------------------------------------------------------- That means \[\large 9*\sqrt[4]{7} -8*\sqrt[3]{7} +4*\sqrt[3]{7} +3*\sqrt[4]{7} \] turns into \[\large 9*x -8*y +4*y +3*x \] hopefully you see how to simplify from here